OpenStudy (anonymous):

On graph paper, graph y = (2/7)x – 2. Explain how to graph the equation using the slope and y-intercept.

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

y=mx+b tells us the slope is m and the y-intercept is (0,b) So can you first identify the slope and the y-intercept ?

OpenStudy (anonymous):

m=27,b=−2 where m is the slope and b is the y-intercept

OpenStudy (freckles):

m=2/7 and b=-2 so the slope is 2/7 and the y-intercept is (0,-2) So your job is to first plot the point (0,-2) this just a point on the y-axis 2 units below the origin

OpenStudy (freckles):

after you plot this point we will use m=2/7 to find one more point slope is rise/run so what m=2/7 tells us is we are going to start from our (0,-2) and count up 2 units from then count right 7 units

OpenStudy (freckles):

and that will be another point on the line then connect the dots and you have your line

OpenStudy (freckles):

so basically if you have (x,y) on the line then (x+7,y+2) will also be on the line

OpenStudy (freckles):

so since you have (0,-2) on the line then what other point can you say is on the line?

OpenStudy (anonymous):

(2,5)??

OpenStudy (freckles):

So we have (x,y) then (x+7,y+2) you have (0,-2) then (0+7,-2+2) right?

OpenStudy (anonymous):

ya

OpenStudy (freckles):

so that means plot both points (0,-2) and (7,0) and connect the dots

OpenStudy (anonymous):

okay :) could you help me on one more?

OpenStudy (freckles):

ok I can try

OpenStudy (anonymous):

On graph paper, graph the line containing the point (–6, 6) and having a slope of m = –3/2. What are the coordinates of a second point that is located on the line?

OpenStudy (freckles):

pretend just like before we know we have a line and we know the point (x,y) is on there and we the slope is b/a what the slope means is we are going to move (x,y) up b units and right a units (assuming a and b are positive) so you will have the point (x+a,y+b) also on the line as you can see if we do \[\frac{(y+b)-y}{(x+a)-x}=\frac{b}{a}\] which was indeed the slope we had for this example if you have point (x,y) on the line then you also have if the slope is b/a then the point (x+a,y+b) is on the line

OpenStudy (freckles):

so you have \[\frac{-3}{2}=\frac{b}{a}\] b=-3 and a=2 so if (-6,6) is on the line then what other point is also on the line?

OpenStudy (anonymous):

(-9,-8)

OpenStudy (freckles):

let see (x,y) -> (x+a,y+b) a=2 and b=-3 x=-6 and y=6 so we have (-6,6) -> (-6+2,6+(-3))

OpenStudy (anonymous):

-8,3)

OpenStudy (freckles):

-6+2 is -4 :p so you have another point on the line is (-4,3)

OpenStudy (freckles):

now I'm not saying these are the only two points on the line it is just another way to find anther point

OpenStudy (freckles):

anyways plot (-6,6) and (-4,3) and connect the dots

OpenStudy (freckles):

and you know you should continue the line past the points after all it is a line lines go in both directions forever and ever

OpenStudy (anonymous):

thank you for your help:D